hek:

i need help pls What is the difference of 823−179? You may use fraction strips to help. 649 689 719 759

hek:

what

AquarytheALPHA:

are you trying to like sub and get the answer?

hek:

i need help

AquarytheALPHA:

ok i can help you

hek:

ok

AquarytheALPHA:

It's A. 649

hek:

what is the difference of 823-179

R3al:

644

hek:

i done got that

AquarytheALPHA:

well is it a fraction? and if it is what do you think it is

hek:

yes it was a fraction

AquarytheALPHA:

0k which answer do you think it is

AquarytheALPHA:

it is either A or B

AquarytheALPHA:

hek you there?

jp:

no he iis doing something he is my little brother

XioGonz:

@hek Have you gotten the answer you were searching for?

jhonyy9:

@hek you need subtract the 179 from 823 can you calcule it ?

hek:

i am here

hek:

no it won't let me

hek:

anybody here

jhonyy9:

823 - 179 how you subtract ?

hek:

so you take 3 from 9

hek:

and 2 from 7

hek:

and 8 from 1

jhonyy9:

no bc. there is from 3 take 9 what not posibil so get a ten from 2 and so will subtract the 9 from 13 so 13 minus 9 = ?

hek:

i done got it

hek:

i am on something hard

jhonyy9:

i wan teach you how subtract step by step - pls collaborate

hek:

i don't no how to do this it ant subtract

jhonyy9:

just pls collaborate and i will teach it to you ok?

hek:

this ant subtracting

paigeeeeeeeeeeee:

omg off topic but i love ur pfp @hek

paigeeeeeeeeeeee:

todoroki </3

hek:

who's

paigeeeeeeeeeeee:

urs

hek:

thank you

paigeeeeeeeeeeee:

yw

jhonyy9:

823- 179 

hek:

i am trying to get help from jhonyy9

jhonyy9:

@hek do you understand ?

hek:

yea

jhonyy9:

this is the way what you need learn it

jhonyy9:

these point up to 2 and 8 this is for you dont forgot these tens what you need get from there to make the subtraction right

jhonyy9:

questions ?

jhonyy9:

have you understood it ?

jhonyy9:

so than try calcule this pls, 122- 99

jhonyy9:

do you remember how i ve said above what is the way ?

jhonyy9:

so how you see nobody dont wan explain it to you

jhonyy9:

do you understand ?

jhonyy9:

122 - 99 ohh hope you know that we begin from right part so from 2 not can subtract 9 so get a ten from second 2 and hence sou subtract 9 from 12 so 12 -9 = ?

jhonyy9:

pls collaborate

jhonyy9:

ok so 12 -9 = 3 continue with second 2 but there is just 1 bc. we get one ten from there to the first 2 so there is now in this way 1 -9 what is not posibil so we get a ten from 1 and so will be 11 - 9 = 2 so in this way we get that 122 - 99 = 23

jhonyy9:

do you understand now ?

jhonyy9:

questions ? so than was my pleasure bye bye

hek:

bye bye